i need to plot at least three points for each graph for 3x-2y=6
Answer by philline_palana(20)

3x-2y=6

if x=0
3(0)-2y=6
-2y=6 >>divide by -2
-2y/-2=6/-2
y=-3
point 1: (0,-3)

if x=1
3(1)-2y=6
3-2y=6 >>transpose 6
3-6-2y=0 >>transpose -2y
3-6=2y
-3=2y >>divide by 2
-3/2=2y/2
y=-3/2
point 2: (1,-3/2)

if y=0
3x-2(0)=6
3x=6 >>divide by 3
3x/3 = 6/3
x=2
point 3: (2,0)